# UnMarker: A Universal Attack on Defensive Watermarking

### v1

@@ 6.2.3. Comparison with SOTA attacks
Overall, among the two baselines, only the DiffusionAttack can remove watermarks without violating visual constraints. Yet, UnMarker remains far superior as it succeeds against semantic watermarks where the DiffusionAttack fails.
7. Discussion
Practicality. UnMarker can be executed offline as it requires no detector interaction, given only the watermarked image. A single attack takes up to $5$ minutes with a $40GB$ A100 GPU. Such machines are available for \$$\$32.77/hr$ on AWS (aws, [n. d.]), making it highly practical.
Existing query-free black-box attacks (see §6.2.3) require advanced denoising pipelines (e.g., VAEs and DMs), whose ability to preserve quality depends on their computational power and the commonalities between their training data and the distributions from which watermarked images are generated and fail when these conditions are not met as we demonstrated for the VAEAttack. While the DiffusionAttack achieves better quality (despite it being inferior to UnMarker’s), the DM (hf, [n. d.]) used in this attack and all the studied schemes were trained on public datasets. In practice, providers possess exclusive high-quality data, which may vary significantly from those used to train the DMs available to the attacker. Thus, the DiffusionAttack may incur significant quality degradation in reality. UnMarker’s data-free nature is agnostic to these nuances.
Mitigations. UnMarker erases the watermark’s footprint. Although it is an adversarial attack, known mitigations (e.g., adversarial training (Tramèr et al., 2017; Hu et al., 2024; An et al., 2024)) are not applicable since, in our unique scenario, UnMarker itself removes an “adversarial” signature (i.e., the watermark) previously injected by the provider. This volatility of watermarks (being adversarial themselves) makes them a weak attestation method: once these fragile signatures are removed, they cannot be restored. This is in contrast to traditional adversarial problems wherein the true class of the sample is retained and may be recovered by discarding the attacker’s minor modifications. Hence, watermark removal cannot evolve into a cat-and-mouse game where providers continue to devise defenses. UnMarker’s persistence comes from its destructive nature and its agnosticism of the target scheme. On the other hand, known adversarial attacks against defensive watermarking (Lukas et al., 2024; Jiang et al., 2023; Saberi et al., 2024) follow “traditional” principles, relying on detector feedback to generate disruptions by exploiting its imperfections, yielding “volatile” adversarial samples lying in uncertainty regions of these systems. This makes them reversible via adversarial defenses, as they may retain the watermark but overshadow it with bounded noise that can be eliminated. Our findings indicate that watermarking is not a viable defense against deepfakes and urge the community to explore alternatives.
@@ 7. Discussion
Ethical Considerations. Misusing GenAI can have severe impacts on society. Our attack does not aim to help generate harmful content but to raise awareness regarding these systems’ risks. Although we are not aware of providers that have already rolled out stable watermarking algorithms, the pace at which they are being adopted, and the attention from entities such as the White House that has secured commitments from providers to deploy these solutions (Heikkilä;, 2023) make it mandatory that they are thoroughly assessed. While we reached out to providers requesting access to their APIs for evaluation, we never heard back. Thus, all our studied systems (including Facebook Research’s StableSignature (Fernandez et al., 2023)) are open-source.
8. Conclusion
@@
We presented a novel attack against defensive watermarking via adversarial optimizations that disrupt the spectra of watermarked images and interfere with the watermarks’ footprints, erasing them without interacting with the scheme. UnMarker is effective against all systems, including semantic schemes that have been assumed robust to such query-free and black-box attacks. Our results expose the risks associated with defensive watermarking and urge providers to invest further efforts in developing more robust deepfake defenses.
@@ Acknowledgments
Appendix A Appendix
A.1. Generative AI
Leading image generators fall under three classes:
Generative Adversarial Networks (GANS). A GAN (Goodfellow et al., 2014) consists of a generator $\mathcal{G}:\mathcal{Z}\longrightarrow\mathcal{X}$, where $\mathcal{Z}$ is a space of latent codes from which $\mathcal{G}$ generates images in $\mathcal{X}$ that represents this desired image distribution, and a discriminator $\mathcal{D}:\mathcal{X}\longrightarrow(0,1]$, that tells apart real from fake images. In practice, $\mathcal{D}$ outputs a probability for the sample being real or fake. The two are simultaneously trained s.t $\mathcal{D}$ learns to accurately discern between real images and those outputted by $\mathcal{G}$, while $\mathcal{G}$ is proactively improved to fool the continuously-enhanced discriminator $\mathcal{D}$ by generating more real-looking data, pushing it to synthesize high-quality images. Afterward, $\mathcal{G}$ can be used to generate images from random latent codes. NVIDIA’s StyleGAN (Karras et al., 2019, 2020; Sauer et al., 2022) is a family of advanced GANs.
[…]
Latent Diffusion Models (LDMs). LDM(Rombach et al., 2022) are DMs that do not operate in the pixel space but rather deploy and encoder $\mathcal{E}$ and a decoder $\mathcal{D}$. $\mathcal{E}$ transforms images into compressed latent representations on which the above diffusion passes are performed. $\mathcal{D}$ then transforms the reconstructed latents back into images. LDMs offer several advantages over DMs such as efficiency and enhanced resolution. A popular LDM is Stability AI’s Stable Diffusion (sta, [n. d.]).
Conditional Generation. All generators can be used in a conditional manner, designed to accept additional information that influences their output, such as text describing the desired content (DAL, [n. d.]; sta, [n. d.]) or the identity of the person whose picture is to be synthesized (Cao et al., 2017).
A.2. Details Regarding The Watermarking Schemes’ Configurations
The watermark lengths used for each of the schemes we study are in Table 4. Recall that we use the official implementations provided by the authors of each scheme, who found these to be optimal. The systems were, therefore, pretrained to operate using these lengths.
Table 4. Watermark lengths. TRW’s watermarks are not binary sequences, and thus, this metric is not applicable
 Scheme Watermark Length (# bits)
 Yu1 128
Yu2 100
HiDDeN 30
PTW 40
StableSignature 48
StegaStamp 100
TRW NA
A.3. UnMarker’s Hyperparameters
Below we present UnMarker’s hyperparameter considerations.
@@ A.3. UnMarker’s Hyperparameters
High-frequency Disruptions (Stage 1— §5.1): We use the Alex-Net-based LPIPS (Zhang et al., 2018) for small images ($128{\times}128$) and the DVL (Czolbe et al., 2020) perceptual loss ($\ell_{p}$) for larger sizes. As we maximize ${\ell_{F}}_{0}$, we want to set the threshold $t_{{}_{\ell_{p}}}$ to the largest value at which the changes still retain similarity and do not lead to artifacts. This $t_{{}_{\ell_{p}}}$ depends on $\ell_{p}$ and the image size (resolution). For small images optimized using the Alex-Net-based loss, this optimal point is $t_{{}_{\ell_{p}}}=5\cdot 10^{-4}$. Since larger images require a different loss (DVL), similar thresholds do not indicate the same degree of resemblance. For $256{\times}256$ images, we find $t_{{}_{\ell_{p}}}=4\cdot 10^{-2}$ best achieves the above goals, while for larger images we have $t_{{}_{\ell_{p}}}=10^{-4}$. Next, we select the initial $c_{{}_{\ell_{p}}}$ that balances the spectral and visual losses. We want this starting value to be significant to assign a large weight to $\ell_{p}$ and ensure visual similarity is not violated, after which it can be decreased to allow the spectral loss to be more dominant while still ensuring this property. This initial setting only affects the time until it converges and must be empirically determined to speed up the process. We found an initial $c_{{}_{\ell_{p}}}=10^{6}$ optimal for small images that use the Alex-Net-based LPIPS. For DVL, this was set at $10^{6}$ when the size is above $256{\times}256$ as these images require the DVL loss to be extremely small ($10^{-4}$), while for $256{\times}256$ images, it can be far smaller ($10^{-2}$).
@@
The $\|\|_{2}$ component prevents unwanted peaks (see §5.1). Its threshold depends on the size and chosen LPIPS. While the LPIPS thresholds guarantee visual similarity to the extent that the images will look similar up to, potentially, such peaks, $\|\|_{2}$’s thresholds must eliminate these singular points when the LPIPS thresholds are met. We empirically selected these $\|\|_{2}$ thresholds: for $128{\times}128$, we have $t_{\ell_{\|\|_{2}}}=10^{-4}$, while $256{\times}256$ images allow $t_{\ell_{\|\|_{2}}}=3\cdot 10^{-5}$, and larger images operate at $t_{\ell_{\|\|_{2}}}=10^{-4}$. We report the normalized thresholds (divided by the image size). In practice, they increase with the size since larger images accommodate stronger perturbations without visible cues. $c_{\ell_{\|\|_{2}}}$ multiplies this $\|\|_{2}$ when its threshold is violated to strongly steer the procedure away from these boundaries. This is not determined through a binary search, as we want it to constantly have the same effect. It should be selected to accomplish this without undoing the majority of the previous iterations’ effects. By tracking the losses, we arrive at $c_{\ell_{\|\|_{2}}}=0.6$.
By monitoring the spectral loss, we can determine when it stops increasing for each value of $c_{{}_{\ell_{p}}}$ and, similarly, the number of steps after which updating $c_{{}_{\ell_{p}}}$ does not result in further improvements. This yields $5000$ iterations for images smaller or larger than $256{\times}256$, while for this size, the loss plateaus after $2000$ iterations. The number of binary search steps is $2$ for sizes $\leq$ $256{\times}256$ and $4$ otherwise.
We use the Adam optimizer with learning rate $2\cdot 10^{-4}$, reduced by $0.5$ when no improvement (above $10^{-4}$) to the spectral loss occurs for $10$ iterations. We clip the gradients at $5\cdot 10^{-3}$ for DVL and $5\cdot 10^{-2}$ for the Alex-Net-based loss to ensure they do not dominate the process and allow the spectral loss to influence the results.
Low-frequency Disruptions (Stage 2— §5.2): Each filter $OF^{t}_{{\mathcal{K}^{*}_{{}_{{}_{\;\;t}}}},x_{w},{\sigma_{c}}_{t}}$ is characterised by two hyperparameters: dimensionality $M_{t}{\times}N_{t}$ of the kernel ${\mathcal{K}^{*}_{{}_{{}_{\;\;t}}}}$ and permissiveness ${\sigma_{c}}_{t}$. We find ${\sigma_{c}}_{t}=5\cdot 10^{-2}$ always optimal. When choosing the kernel sizes and number of filters $T$, resources can become a bottleneck as we have a kernel per pixel, making each filter scale as the number of pixels multiplied by the kernel size. We found $T=9$ with different kernel sizes defeats all studied systems. This architecture was chosen (semi-randomly) to target various potential geometries of the watermark under our computational limitations ($40$ GB A100 GPU) but can be further optimized (additional layers) with better hardware. However, our findings indicate that this may not be necessary. The chosen sizes (in sequential order) are (21, 5),215(21,\;5),( 21 , 5 ) , (5, 5),55(5,\;5),( 5 , 5 ) , (17, 33),(7, 7),(47, 5),(33, 17),(17, 17),(5, 5)1733774753317171755(17,\;33),\;(7,\;7),\;(47,\;5),\;(33,\;17),\;(17,\;17),\;(5,\;5)( 17 , 33 ) , ( 7 , 7 ) , ( 47 , 5 ) , ( 33 , 17 ) , ( 17 , 17 ) , ( 5 , 5 ) and (3, 3)33(3,\;3)( 3 , 3 ).
The spectral loss (see eq. (6)) consists of MPL and FRL. While we maximize both, FRL is effectively a regularization term supporting MPL— see §5.2.4. Hence, we focus on MPL, deeming a sample to be optimal if it maximizes this loss. Yet, $\lambda_{{}_{\tiny\mbox{$FRL$}}}$ must account for FRL, ensuring the outputs meet the conditions it enforces. $\lambda_{{}_{\tiny\mbox{$FRL$}}}=5$ was found optimal, but it can be adjusted if the two components do not meet expectations (see §5.2.4). We found the $\|\|_{2}$ term’s coefficient $\lambda_{{}_{\tiny\mbox{$\|\|_{2}$}}}=2.5\cdot 10^{-4}$ ideal but it can be calibrated similarly.
@@
Table 5. Performance of the schemes against basic manipulations. The first row shows the results in the absence of any manipulations. Numbers in parentheses are the detection thresholds, while Detect is the portion of images meeting them. For attacks with multiple configurations, lower rows correspond to stronger distortions. * denotes overly-destructive configurations.
    Yu1 (61%) Yu2 (63%) HiDDeN (73%) PTW (70%) StableSignature (69%) StegaStamp (63%) TRW (0.0141)
@@
For low-frequencies, DVL is universally applicable (see §5.1). The remaining unknowns were chosen based on similar considerations to those for our high-frequency stage. As low-frequency perturbations do not risk introducing unwanted peaks, these parameters can be selected uniformly with little regard to the image size. The exception is the perceptual threshold $t_{{}_{\ell_{\tiny\mbox{${DVL}$}}}}$, as smaller images ($256{\times}256$ and below) are more sensitive to structural changes (as opposed to high-frequency perturbations that can be masked by these images’ low resolution), exhibiting (minor) visible differences for $t_{{}_{\ell_{\tiny\mbox{${DVL}$}}}}>10^{-3}$, while larger images can withstand three times this change. With an initial $c_{{}_{\ell_{\tiny\mbox{${DVL}$}}}}=10^{6}$, a single step running for $500$ iterations is sufficient. We use Adam with a $10^{-2}$ learning rate for the filters and $2\cdot 10^{-3}$ for the modifier $\delta$, decayed by $0.99$ if MPL fails improve by $5\cdot 10^{-2}$ for $10$ iterations. As DVL’s impact is weakened due to the structural changes this process induces, we allow its gradients to be more dominant, clipped at $1.0$.
@@
Filters operate on all image channels identically. As the reason behind including a direct modifier $\delta$ here is the filters’ ability to smoothen these otherwise high-frequency changes, it must affect all channels similarly for them to achieve this outcome. Thus, we use a (gray-scale) modifier $\delta$ that applies the same changes to all channels. Finally, watermarks exist in the spectral amplitudes rather than phases. While manipulating amplitudes impacts the texture and, therefore, results in less critical changes, phase modifications are more visible (see §4.1.1). Hence, we should ideally restrict our filters to manipulate these magnitudes. The frequency response of a filter does not change the phase if and only if it is even-symmetric (identical to the reflections of each of its quadrants across both axes). While this holds for basic (convolutional) filters, ours will converge to similar kernels for neighboring pixels that are of equal visual importance and can be altered similarly, indicating that each region will behave as if it were “convolved with a basic kernel”. Thus, our filters should possess this even-symmetry property. Accordingly, instead of optimizing all kernel weights, we focus only on the first quadrant of each filter and then reflect it across both axes.
A.4. Baseline Results Against Naive Image Manipulations
Most systems are robust to all practical manipulations. Interestingly, Yu2 breaks completely under cropping. This result does not contradict the original paper (Yu et al., 2021); The authors watermark several generators yet only evaluate a single ProGAN (Karras et al., 2017) architecture under manipulations, whereas we chose their advanced StyleGAN2 (Karras et al., 2020) model. Overall, our observations reveal that, unlike all other schemes, Yu2 is highly dependent on spatial attributes, making it volatile to manipulations such as cropping. However, it remains robust to all other modifications. Several other schemes (Yu1, HiDDeN and TRW) exhibit a considerable drop in the average bit accuracies (inverse distance) after cropping, despite their overall detection rates not being affected. This is not surprising, as the spectral modifications for watermark injection must be spread throughout the image (see §4.1) to minimize visual cues, implying that a portion may be allocated to the background. Yet, they remain robust by restricting most of these changes to central regions.
Yu1, StegaStamp and TRW resist all manipulations, while StableSignature can withstand realistic, non-obtrusive modifications, making it also robust for practical purposes. PTW is slightly less powerful since it loses some of its efficacy under even non-destructive operations, such as Quantization with a factor of $10$ and JPEG with quality $80$ that decrease its performance to $46\%$ and $76\%$, respectively. HiDDeN exhibits similar behavior, showing resistance to all realistic manipulations, aside from JPEG that brings its detection to $43\%$ even when used with a reasonably good quality of $110$ or Blurring with $std.=1.0$ that lowers its detection score to $5.0\%$.
It is not unnatural to see that some non-semantic systems (e.g., PTW, HiDDeN and StableSignature) begin to break under spectral manipulations such as JPEG that has been shown effective against heuristics-based watermarks (see §2) operating in the spectrum as well, Blurring or Quantization. As explained in §4.1, constructing spectral non-semantic watermarks should be done in an input-specific manner that is also persistent, which is a challenging task that will be restricted to volatile watermarks unless the scheme is sufficiently advanced. This explains why the more sophisticated StableSignature and PTW outperform HiDDeN in this department. Yet, allowing for more severe spectral manipulations by applying stronger (Guided) Blurring breaks StableSignature as well, although this experiment is only included to demonstrate this phenomenon that attests to the spectral nature of these watermarks and does not make for a valid attack, as explained in §6.1. Similarly, Quantization can cause sharp pixel shifts and affect the signal’s smoothness, introducing high-frequency noise, explaining PTW’s vulnerability.
For all naive manipulations, our findings are comparable to those attained by the authors of the schemes that evaluate the same attacks with similar parameters. The only exceptions are Gaussian Blurring with std.=1.0 against HiDDeN and JPEG with quality=80 against PTW, for which we observed a non-negligible drop in the bit accuracies compared to these original works. We hypothesize that this is due to us using a different kernel size for the Gaussian filter (the authors of HiDDeN do not report this parameter) and a different implementation of the JPEG algorithm compared to PTW.

### v2

@@ 6.2.3 Comparison with SOTA attacks
Among the SOTA attacks, only the DiffusionAttack can remove watermarks without violating visual constraints. Yet, UnMarker is superior as it succeeds against semantic watermarks where the DiffusionAttack fails.
7 Discussion
Practicality. UnMarker is executed offline as it requires no detector interaction and takes up to $5$ minutes with a $40GB$ A100 GPU. Such machines are available for \$$\$32.77/hr$ on AWS. It requires no data, freeing it from the dependence on denoising pipelines (e.g., DMs). Regeneration attacks are affected by these systems’ fidelity and the commonalities between their training data and the watermarked images, as we showed for the VAEAttack. While the DiffusionAttack achieves better quality (but worse than UnMarker’s), its DM and all studied schemes were trained on public data. Providers possess exclusive data that differ from those used to train the attacker’s DM, making it likely to impact quality.
Mitigations. In the context of defensive watermarking, mitigations should allow recovery from removal attacks, enabling providers to continue using the scheme. Adversarial training (AT[61] is a popular mitigation in ML, wherein the model (detector) is trained on malicious samples from known attacks so that it learns to classify them correctly. Mitigations can also target the model’s inputs: the provider applies a pre-processing stage (e.g., purification [62]) to eliminate disruptions and only then invokes the model (detector). Mitigations could exist for adversarial attacks as they typically retain the sample’s true class but rely on model feedback to exploit its imperfections, yielding “volatile” perturbations in its uncertainty regions that cause misclassification. This strategy has also been the focus of previous work on watermark removal [32, 30, 29]. Yet, if the watermark is actually removed, recovery would be impossible. As we claim that by simply disrupting the carrier, UnMarker removes the watermark, mitigations should fail.
To test this, we consider a SOTA mitigation— diffusion purification [62]. Purification inspires the DiffusionAttack (see §6.2.3), operating similarly: It adds noise, diluting adversarial perturbations, and uses a DM for reconstructing a pure (attack-free) sample. As this applies as a certifiable attack on non-semantic watermarks, these schemes cannot use this mitigation. Yet, these watermarks are of little importance since these proofs for their removal undermine their practicality. As we aim to investigate the validity of defensive watermarking, we focus on the more reliable semantic schemes that meet the standards for utilizing purification as a defense. Still, UnMarker’s stages against both classes hinge on similar principles, making the findings applicable to defenses for non-semantic schemes if implemented.
We attack $100$ samples for each scheme using UnMarker (CHL). Then, we perform purification and invoke the detector. Given the resistance of semantic schemes to purification (the DiffusionAttack) and the effectiveness of this mitigation against adversarial examples, if UnMarker does not remove the watermark but overshadows it with volatile changes, purification would eliminate UnMarker and restore the watermark. Yet, the results verify UnMarker’s destructive nature that makes recovery impossible. For purification with $t\!\!=\!\!0.1$ (see §6.2.3), StegaStamp’s detection drops to $15\%$ and TRW’s to $16\%$. A larger $t$ leads to worse detection. Ironically, the defense degrades performance. While this is partly because diffusion may slightly decrease the bit accuracy (inverse distance), we apply it at half of its original strength (see §6.2.3), which allowed for $\sim\!\!100\%$ detection. The reason for this decline is UnMarker’s destruction, impairing the DM’s reconstruction ability. This raises the possibility of combining UnMarker with the DiffusionAttack as an even stronger attack, which we leave to future work due to the potential increased quality degradation. Evaluating AT is prohibitive due to time limitations, as it requires numerous adversarial examples, and a single UnMarker attack takes several minutes. Yet, AT is only successful if the attack sample retains the watermark, which is not the case for UnMarker, as proven. As purification adheres to similar principles, its ineffectiveness implies the same for AT.
Outlook. A natural question is whether watermarking can ever be made reliable. Particularly, can we devise schemes that do not conform to our spectral analysis s.t. UnMarker, or similar attacks, would fail? The answer is likely no.
In §4, we establish that a universal carrier that encodes the watermark is crucial for any scheme due to computational considerations. Given the nature of visual content, the assertion that this carrier must be in the spatial structure or spectrally constructed is also invariant. Then, we explain how spatial dependencies lead to fragile watermarks, as is demonstrated in §4.1. Thus, robust schemes must embed watermarks in the collective spectral amplitudes. Non-semantic schemes are unreliable and, despite their slightly demanding nature, regeneration attacks certifiably erase them, restricting strong watermarks to lower frequencies (semantic). While these are of larger magnitudes, they must remain minimally visible to retain quality and since obvious traces are removable even via naive editing. Existing semantic schemes push these demands to the extreme, leaving somewhat visible footprints (StegaStamp) or changing the content and limiting the space of watermarked images (TRW). Since UnMarker defeats them, future variants will have to further influence the content. Due to the considerable modifications the SOTA methods incur, such schemes will likely be impractical, reducing the quality and variety of watermarkable content and leaving visible traces. Combined with the inefficacy of mitigations, defensive watermarking becomes invalid, and future work should explore alternatives.
Limitations & Future Work. 1) We focus on practical (current) systems applicable to various images. Future schemes might be tailored for specific content (e.g., adding brand-specific glasses to face images), but such GenAI services will lack diversity. To meet users’ needs, providers will also offer less restrictive tools, and attackers can use them instead, rendering content-specific watermarking useless. To reduce visibility and resist naive attacks, they must also employ spectral manipulations (for reasons similar to §4.1.1’s), leaving UnMarker effective. Yet, without requiring universal applicability, the spectral footprint may target certain regions or objects only, not the collective spectrum. This necessitates augmenting UnMarker with content-based knowledge to identify optimal disruption regions, which could be studied should such schemes become prevalent. 2) UnMarker relies on accurate visual losses that preserve essential similarity cues while maximizing perturbations to attain the best removal rates, and future work may explore alternatives superior to ours (see §5.1). 3) While we consider visual content, our analysis translates to other domains (e.g., audio), but UnMarker requires domain-based adaptations that can be interesting to investigate. 4) As watermarks bear resemblance to adversarial examples (AEs), UnMarker could potentially defend against AEs. Still, as AEs do not require a universal carrier, they may target other channels (aside from collective spectra), mandating further improvements.
@@ 7 Discussion
Ethical Considerations. Misusing GenAI has severe impacts on society. We aim not to help generate harmful content but raise awareness regarding these systems’ risks. Although we are not aware of providers that already released stable watermarking schemes, their adoption pace and attention from dominant entities (see §1) make assessments mandatory. We requested access to providers’ APIs for evaluation but never heard back. Thus, all studied systems are open-source.
8 Conclusion
@@
We presented a novel attack against defensive watermarking via adversarial optimizations that disrupt the spectra of watermarked images and interfere with the watermarks’ footprints, erasing them without interacting with the scheme. UnMarker is effective against all systems, including semantic schemes that have been assumed robust to such query-free and black-box attacks. Our results show that defensive watermarking is not a viable countermeasure against deepfakes and that more robust defenses are needed.
@@ Acknowledgments
Appendix A UnMarker’s Hyperparameters
@@
High-frequency Disruptions (Stage 1— §5.1): We use Alex [51] for small images ($128{\times}128$) and DVL [52] for larger sizes as the perceptual loss ($\ell_{p}$). As we maximize DFL, we set $t_{{\ell_{p}}}$ to the largest value at which similarity and quality are retained. This depends on $\ell_{p}$ and the image size. For small images optimized using Alex, this optimal point is $t_{{{\text{\tiny\mbox{${Alex}$}}}}}\!\!\!=\!\!5\!\cdot\!10^{-4}$. Since larger images require a different loss (DVL), similar thresholds do not indicate the same degree of resemblance. For $256{\times}256$ images, we find $t_{{{\text{\tiny\mbox{${DVL}$}}}}}\!\!=\!4\!\cdot\!10^{-2}$ best achieves the above goals, while for larger images we have $t_{{{\text{\tiny\mbox{${DVL}$}}}}}\!\!=\!\!10^{-4}$. Next, we select the initial $c_{{\ell_{p}}}$ that balances the spectral and visual losses. We want this starting value to be significant to assign a large weight to $\ell_{p}$ and ensure visual similarity is not violated, after which it can be decreased to allow the spectral loss to be more dominant while still ensuring this property. This initial setting only affects the time until it converges and must be empirically determined to speed up the process. We found an initial $c_{{{\text{\tiny\mbox{${Alex}$}}}}}\!\!=\!\!10^{6}$ optimal for small images. For DVL, this was set at $c_{{{\text{\tiny\mbox{${DVL}$}}}}}\!\!=\!\!10^{6}$ when the size is above $256{\times}256$ as these images require the DVL loss to be extremely small ($t_{{{\text{\tiny\mbox{${DVL}$}}}}}\!\!=\!\!10^{-4}$), while for $256{\times}256$ images, it can be far smaller ($c_{{{\text{\tiny\mbox{${DVL}$}}}}}\!\!=\!\!10^{-2}$).
@@
The $\|\|_{2}$ component prevents unwanted peaks (see §5.1). Its threshold depends on the size and chosen LPIPS. While the LPIPS thresholds guarantee visual similarity to the extent that the images will look similar up to, potentially, such peaks, $\|\|_{2}$’s thresholds must eliminate these singular points when the LPIPS thresholds are met. We empirically selected these $\|\|_{2}$ thresholds: for $128{\times}128$, we have $t_{{\|\|_{2}}}=10^{-4}$, while $256{\times}256$ images allow $t_{{\|\|_{2}}}=3\cdot 10^{-5}$, and larger images operate at $t_{{\|\|_{2}}}=10^{-4}$. We report the normalized thresholds (divided by the image size). In practice, they increase with the size since larger images accommodate stronger perturbations without visible cues. $c_{{\|\|_{2}}}$ multiplies this $\|\|_{2}$ when its threshold is violated to strongly steer the procedure away from these boundaries. This is not determined through a binary search, as we want it to have the same effect constantly, without undoing the previous iterations’ effects. By tracking the losses, we arrive at $c_{{\|\|_{2}}}=0.6$.
TABLE IV: Performance of watermarking schemes against basic manipulations.
Note: The first row shows the results in the absence of any manipulations. Numbers in parentheses are the detection thresholds, while Detect is the portion of
@@
the images meeting them. For attacks with multiple configurations, lower rows correspond to stronger distortions. * denotes overly-destructive configurations.
    Yu1 (61%) Yu2 (63%) HiDDeN (73%) PTW (70%) StableSig. (69%) StegaStamp (63%) TRW (0.0141)
@@
By monitoring the spectral loss, we can determine when it stops increasing for each value of $c_{{\ell_{p}}}$ and the number of steps after which updating $c_{{\ell_{p}}}$ does not result in improvements. This yields $5000$ iterations for images smaller or larger than $256{\times}256$, and $2000$ otherwise. The number of binary steps is $2$ for sizes $\leq$ $256{\times}256$ and $4$ otherwise. The optimizer is Adam with learning rate $2\cdot 10^{-4}$, reduced by $0.5$ when no improvement (above $10^{-4}$) occurs for $10$ iterations. We clip gradients at $5\cdot 10^{-3}$ for DVL and $5\cdot 10^{-2}$ for Alex to ensure they do not dominate the process.
Low-frequency Disruptions (Stage 2— §5.2): Each filter is characterised by the dimensionality $M_{t}{\times}N_{t}$ of ${\mathcal{K}^{*}\!\!\!_{{}_{t}}}$ and permissiveness $\!\!{\sigma\!_{c_{{}_{t}}}}$. We find ${\sigma\!_{c_{{}_{t}}}}\!\!=\!\!5\cdot 10^{-2}$ optimal. When choosing the kernel sizes and number of filters $T$, resources may be prohibitive as we have a kernel per pixel, making a filter scale as the number of pixels multiplied by its kernel size. We found $T\!\!\!=\!\!9$ with different sizes defeats all studied systems. This architecture was chosen (semi-randomly) to target various potential geometries of the watermark under our computational limitations ($40$GB A100 GPU) but can be extended with better hardware. Yet, our findings indicate that this may not be necessary. The ordered sizes are $(21,5),\!$ $(5,5),(17,33),(7,7),(47,5),(33,17),(17,17),(5,5)\!\!\!$ and $(3,3)$. Our spectral loss consists of MPL and FRL. While we maximize both, FRL is a supporting regularization term— see §5.2.4. Hence, we deem a sample optimal if it maximizes MPL. Yet, $\lambda_{{\text{\tiny\mbox{$FRL$}}}}$ must account for FRL, ensuring the outputs meet the conditions it enforces. $\lambda_{{\text{\tiny\mbox{$FRL$}}}}\!\!=\!\!5$ was found optimal, but it can be adjusted if the two components do not meet expectations (see §5.2.4). We found $\lambda_{{\|\|_{2}}}\!\!=\!\!2.5\cdot 10^{-4}$ ideal but it can be calibrated similarly.
@@
We universally use DVL (see §5.1). As low-frequency perturbations do not introduce sporadic peaks, the remaining parameters can be selected uniformly with little regard to the image size. The exception is the threshold $t_{{{\text{\tiny\mbox{${DVL}$}}}}}$, as smaller images ($\leq 256{\times}256$) are more sensitive to structural changes (unlike high-frequency perturbations that these images’ low resolution masks), exhibiting visible differences for $t_{{{\text{\tiny\mbox{${DVL}$}}}}}>10^{-3}$, while larger images can withstand 3×3$\times$3 × this change. With an initial $c_{{{\text{\tiny\mbox{${DVL}$}}}}}=10^{6}$, a single step running for $500$ iterations is sufficient. We use Adam with a $10^{-2}$ learning rate for the filters and $2\cdot 10^{-3}$ for $\delta$, decayed by $0.99$ if MPL fails improve by $5\cdot 10^{-2}$ for $10$ iterations. As DVL’s impact is weakened due to this procedure’s structural nature, we allow its gradients to be more dominant, clipped at $1.0$.
@@
Filters operate on all image channels identically. Hence, as the reason behind including the modifier $\delta$ is the filters’ ability to smoothen its otherwise high-frequency changes, $\delta$ must also affect all channels similarly. Thus, we use a gray-scale modifier $\delta$ that applies the same changes to all channels. While manipulating spectral amplitudes impacts the texture and results in less critical changes, phase modifications are more visible (see §4.1.1). Hence, we should ideally restrict our filters to manipulate magnitudes. The frequency response of a filter does not change the phase iff it is even-symmetric (identical to the reflections of each of its quadrants across both axes). While this holds for basic (convolutional) filters, ours will converge to similar kernels for neighboring pixels that are of equal visual importance and can be altered similarly, indicating that each region will behave as if it were “convolved with a basic kernel”. Thus, our filters should possess this symmetry. Accordingly, instead of optimizing all weights, we focus only on the first quadrant of each kernel and reflect it across both axes.
Appendix B Watermarking Schemes’ Configuration
The (author-reported) watermark lengths are in Table V.
TABLE V: Watermark lengths.
Note: TRW’s watermarks are not binary; thus, this metric is not applicable.
 Scheme Length (# bits)
 Yu1 128
Yu2 100
HiDDeN 30
PTW 40
StableSignature 48
StegaStamp 100
TRW NA
Appendix C Baseline Results Against Image Manipulations